OpenStudy (anonymous):

Solve the following system of equations by linear combination: 3x + 2y = 7 y = x – 4 (3, –1) (2, 6) (–1, 3) (6, 2)

OpenStudy (nikato):

Use substitution

OpenStudy (anonymous):

Substitution?

OpenStudy (nikato):

Yes, like what u did with the x in the last question

OpenStudy (anonymous):

This one isn't hard since they have already solved for y. Now plug in x-4 for y in the first equation. You get ... 3x + 2(x-4)=7. Now expand the paranthesis by distributing the 2 and we get 3x + 2x-8=7. Add like terms to get 5x-8=7 we're solving for x so move 8 to the other side and it becomes 5x=7+8 and that becomes 5x=15. divide both sides by 5 and get x = ?. Plug x into the second equation to get y...so y= ? - 4

OpenStudy (anonymous):

what do i substitute?

OpenStudy (nikato):

The value of y into the first equation

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

substitute x -4 for y in the first equation

OpenStudy (anonymous):

3x + 2-4=7? @Tay003

OpenStudy (anonymous):

3x + 2(x-4) =7

OpenStudy (anonymous):

so whats the answer 0-0?

OpenStudy (anonymous):

im confused, are there more steps? @Tay003

OpenStudy (anonymous):

if you distribute you get 3x +2x -8=7 Always combine like terms so it 5x = 15 divide by 5 and you get the x value

OpenStudy (anonymous):

so 15/5?

OpenStudy (anonymous):

thats three so the first choice?

OpenStudy (anonymous):

yup :)